2. Let X have the pdf Ix(x) = .. ti, 0 < x < 2. Find the pdf of Y X2/2 and P(0 <Y < 1).
Answer #1

here X =-/+√2Y

for |dx/dy| =|(d/dy)*√2Y| =1/(√(2*y))

hence pdf of Y :f(y) =|dx/dy|*(fx(g(y)) =1/(√(2*y))*((2Y)3/2/8+1/4)

f(y)=1/(√(2*y))*((√2Y)3/2/8+1/4) for 0 <y<2

P(0<y<1) =P(0<X2/2 <1) =P(0<X<√2) =\int_{0}^{\sqrt{2}} f(x) dx =\int_{0}^{\sqrt{2}}(x3/8+1/4) dx=(x4/32+x/4)|√20 =(1/8+√2/4)